OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x+y=7 9x+( )y=( )

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Read the definition of coincident here: https://www.mathsisfun.com/definitions/coincident.html

OpenStudy (anonymous):

please help me and @M_palomares and @iamabarbiegirl and @IMANINJA127

OpenStudy (anonymous):

i see

Directrix (directrix):

Look above at the link so that you will know what coincident lines are.

Directrix (directrix):

Then, you will have an idea about the missing coefficients.

OpenStudy (anonymous):

oow

Directrix (directrix):

Look at the two equations in the post. For the x coefficients, what do you do to 3 to get 9. Multiply it by what?

OpenStudy (anonymous):

@ellenalexis11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3

Directrix (directrix):

I am trying to help you and you are tagging others to help you without even giving me a chance.

OpenStudy (anonymous):

right you have to multiply by 3

OpenStudy (anonymous):

it wont happen again sorry

Directrix (directrix):

Look at the coefficients of the y terms in the given equations. The first equation has a y term with an understood coefficient of 1. So, multiply the 1 by what to get the missing y coefficient of equation 2?

OpenStudy (anonymous):

:/ im stuck here i need your help

Directrix (directrix):

What did you multiply by on the x terms - same number to multiply 1 by.

OpenStudy (anonymous):

and after that

OpenStudy (anonymous):

do i do it again

Directrix (directrix):

3x + y = 7 9x+ ( 3 ) y = ( )

Directrix (directrix):

And, what goes in the last "blank?"

OpenStudy (anonymous):

4 right

Directrix (directrix):

No.

Directrix (directrix):

>>>> right you have to multiply by 3

Directrix (directrix):

Every member of this first equation is multiplied by 3. 3x+y=7

Directrix (directrix):

What is 7 times 3 =

OpenStudy (anonymous):

21

Directrix (directrix):

Right

Directrix (directrix):

3x + y = 7 9x+ 3 y = 21

OpenStudy (anonymous):

so it will look like this 9x+3y=21

OpenStudy (anonymous):

ok thanks

Directrix (directrix):

The two equations have the same graph.

Directrix (directrix):

It looks like that but the entry required on the computer answer may be must those coefficients. You'll have to look and determine that.

OpenStudy (anonymous):

ok